. 1. IF A + B + C = 0 then the quadratic equation 3ax 2 + 2bx + c = 0 has, (a) At least one root in (0,1) (b) One root in (2,3) and other in (-2,-1) (c) imaginary roots (d) None of these

.    


 


1.     IF A + B + C = 0 then the quadratic equation 3ax2 + 2bx + c = 0 has,< how to solve>


(a)    At least one root in (0,1)


(b)    One root in (2,3) and other in (-2,-1)


(c)     imaginary roots


(d)    None of these

Grade:12

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Please find the answer to your question below
Let f(x)=ax3+bx2+cx+d
then f(0)=0 and f(1)=a+b+c=0
also f is differentiable on R
By rolles theorem there exists a root(p) in (0,1) such that f1(p)=0
p=(0,1) such that 3ap2+2bp+c=0